OpenStudy (anonymous):

help please...A 18.01 g sample of N2 reacts with 3.02g of H2 to form NH3. what mass of ammonia is formed?

OpenStudy (aaronq):

First, write an equation and balance it. Then, find limiting reactant (convert to moles and divide each amount of moles by it's stoichiometric coefficient in the reaction) after you've identified the limiting reactant, use those amount of moles to find moles of ammonia. Use the stoichiometric coefficients to build a ratio, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) from here, you can solve for moles then convert to grams. To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.

OpenStudy (anonymous):

thank you

OpenStudy (aaronq):

no problem